# Uncovering flow and deformation regimes in the coupled fluid-solid vestibular system

### v1

@@ 2.4 Scalings and non-dimensionalization
$\hat{r}=ar,\quad\hat{s}=Rs,\quad\hat{t}=\mathcal{T}t,\quad\hat{w}=\frac{R\dot{\Omega}_{0}a^{2}}{\nu}w,$ (9)
$\hat{p}=\frac{BR\Omega_{0}}{a^{2}\nu}p,\quad\hat{\eta}=\frac{a^{2}R\Omega_{0}}{\nu}\eta,\quad\frac{d\hat{\Omega}}{d\hat{t}}=\dot{\Omega}_{0}f(t),\quad\hat{\Omega}=\Omega_{0}F(t).$ (10)
(‘Unhatted’ variables are therefore dimensionless counterparts of the corresponding hatted variables.) Here $\mathcal{T}$ is the timescale of variation of the forcing so that $\dot{\Omega}_{0}=\Omega_{0}/\mathcal{T}$. The velocity scale $\mathcal{U}=R\dot{\Omega}_{0}a^{2}/\nu$ is chosen to balance the viscous forces with the Euler force. The pressure scale is chosen to balance the bending term and the forcing term in the plate equation (6). Note that $F(t)$ is the dimensionless angular velocity and $f(t)$ is the dimensionless angular acceleration so that by definition $f(t)=\dot{F}(t)$.
@@ 2.4.1 Dimensionless equations
Following the rescaling of equations (4)–(7), the dimensionless continuity equation (4) reads
@@
$\begin{split}\epsilon&\mathrm{St}\frac{\partial u}{\partial t}+\epsilon^{3}\mathrm{Re}\left(u\frac{\partial u}{\partial r}+\frac{v}{r}\frac{\partial u}{\partial\theta}+\frac{w}{h}\frac{\partial u}{\partial s}-\frac{v^{2}}{r}-\frac{1}{\epsilon}\frac{w^{2}}{h}\cos\theta\right)\\ &-\dot{\Omega}_{0}\mathcal{T}^{2}\mathrm{St}F(t)w\cos\theta-\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}h\cos\theta\\ &=-\kappa\frac{1}{\epsilon}\frac{\partial p}{\partial r}+\frac{\epsilon}{rh}\left[\frac{r}{h}\frac{\partial}{\partial s}\left(\epsilon^{2}\frac{\partial u}{\partial s}-\frac{\partial}{\partial r}(hw)\right)-\frac{1}{r}\frac{\partial}{\partial\theta}\left(\frac{h}{r}\frac{\partial}{\partial r}(r\hat{v})-\frac{h}{r}\frac{\partial\hat{u}}{\partial\theta}\right)\right],\end{split}$ (12a)
$\begin{split}\epsilon&\operatorname{St}\frac{\partial v}{\partial t}+\epsilon^{2}\operatorname{Re}\left(\epsilon u\frac{\partial v}{\partial r}+\epsilon\frac{v}{r}\frac{\partial v}{\partial\theta}+\epsilon\frac{w}{h}\frac{\partial v}{\partial s}+\epsilon\frac{uv}{r}+\frac{w^{2}}{h}\sin\theta\right)\\ &+\dot{\Omega}_{0}\mathcal{T}^{2}\mathrm{St}F(t)w\sin\theta+\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}h\sin\theta\\ &=-\kappa\frac{1}{\epsilon}\frac{1}{r}\frac{\partial p}{\partial\theta}+\frac{\epsilon}{h}\frac{\partial}{\partial r}\left[\frac{h}{r}\left(\frac{\partial}{\partial r}(rv)-\frac{\partial u}{\partial\theta}\right)\right]-\frac{\epsilon}{rh^{2}}\left[\frac{\partial}{\partial\theta}\left(h\frac{\partial w}{\partial s}\right)-\epsilon^{2}r\frac{\partial^{2}v}{\partial s^{2}}\right],\end{split}$ (12b)
$\begin{split}\mathrm{St}&\frac{\partial w}{\partial t}+f(t)(1+\epsilon r\cos\theta)+\epsilon\mathrm{St}\mathcal{T}^{2}\dot{\Omega}_{0}F(t)(u\cos\theta-v\sin\theta)\\ +&\epsilon^{2}\operatorname{Re}\left(u\frac{\partial w}{\partial r}+\frac{v}{r}\frac{\partial w}{\partial\theta}+\frac{w}{h}\frac{\partial w}{\partial s}+\epsilon\frac{uw}{h}\cos\theta-\epsilon\frac{vw}{h}\sin\theta\right)\\ &=-\kappa\frac{1}{h}\frac{\partial p}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left[\frac{r}{h}\left(\frac{\partial}{\partial r}(hw)-\epsilon^{2}\frac{\partial u}{\partial s}\right)\right]+\frac{1}{r^{2}}\frac{\partial}{\partial\theta}\left[\frac{1}{h}\frac{\partial}{\partial\theta}(hw)-\epsilon^{2}\frac{r}{h}\frac{\partial v}{\partial s}\right].\end{split}$ (12c)
@@
$\nabla^{4}\eta-b^{-1}\nabla^{2}\eta+\varrho\left(\operatorname{St}\frac{\partial^{2}\eta}{\partial t^{2}}+\frac{d\Omega}{dt}\right)=\Delta p.$ (13)
The non-dimensionalization procedure introduces several dimensionless parameters. We shall see that the most important of these are the stiffness
$\kappa=\frac{B\mathcal{T}}{a^{2}R\mu},$ (14)
@@
which measures the time scale of vorticity diffusion across the channel width, $a^{2}/\nu$, to the time scale of motion, $\mathcal{T}$. (Note that this version of the Stokes number arises in Stokes’ second problem and is sometimes replaced by the Womersley number, $\mathrm{Wm}=\mathrm{St}^{1/2}$.) We also introduce the Reynolds number of the flow, $\mathrm{Re}=\rho\mathcal{U}R/\mu$, as well as the cupula’s inertia, $\varrho=\rho_{s}a^{2}t_{h}\nu/(B\mathcal{T})$ and inverse bendability $b=B/(Ta^{2})$ (Davidovitch et al., 2011). To understand the relative size (and hence importance) of these parameters, we now discuss characteristic parameter values and typical sizes of dimensionless parameters next.
@@ 2.4.2 Parameter values
First, we consider the geometrical parameters, taken from Daocai et al. (2014): $R\approx 3.2\times 10^{-3}$ m, and $a\approx 1.6\times 10^{-4}$ m, so that the aspect ratio is $\epsilon\sim 0.05$. The cupula’s thickness is usually quoted as $t_{h}\approx 400\,\mu$m.
@@
The most challenging parameters to identify are $B$ and $T$. We can infer the value of the bending stiffness from the results of Selva et al. (2009), who suggest an extremely low value of the Young’s modulus, $E\sim 5$ Pa. Other authors quote simply an estimate for the bending stiffness, $B\sim 10^{-10}\mathrm{~{}N\,m}$ (Rabbitt & Damiano, 1992). Since the cupula is difficult to access and measure, determining the pretension $T$ is significantly more challenging and typically requires detailed experiments. However, incorporating $T>0$ adds no additional complexity to our method of solution, and so we retain it for completeness; the effect of this tension is characterized through the inverse bendability parameter, $b=B/(Ta^{2})$ Davidovitch et al. (2011). We refer to the limit $b\to\infty$ as the plate regime and $b\to 0$ as the membrane regime.
@@
In terms of the motion, usual ranges of operation for humans are $\dot{\Omega}_{0}\sim 1$ s-2 and $\mathcal{T}$ can be anywhere between 0.01 and 10 seconds. We are now in a position to make informed estimates of the sizes of the dimensionless groups appearing in (12)-(13). Both the Stokes number and the relative stiffness $\kappa$ depend on the forcing time scale $\mathcal{T}$. In particular, $\mathrm{St}\ll 1$ for $\mathcal{T}>1$, but the Stokes number may be large for faster movements. Similarly, $\kappa$ can acquire a large range of values. More usefully, the plate’s inertia $\varrho<10^{-3}$ for all physical values of $\mathcal{T}$ and hence may be safely neglected if required. Similarly, the reduced Reynolds number, $\epsilon^{2}\mathrm{Re}\approx 10^{-3}$ and fluid inertia may be neglected. Given the uncertainty in the modulus of the cupula itself, we note only that the inverse bendability, $b=B/(Ta^{2})$, is expected to be large.
@@
The size of the dimensionless groups discussed above will inform our choices when neglecting terms in the governing equations (11)-(12). In particular, we will exploit the smallness of $\epsilon$ to neglect terms of order $\epsilon^{2}\mathrm{Re}$, and the smallness of $\varrho$ to neglect inertial terms in the plate equation (13). The last simplification is particularly useful as it will allow us to simplify spatio-temporal PDEs into two sets of ODEs.
@@
We now turn to consider in more detail the behaviour of the model for the typical parameter values already discussed. Given the broad range of values that may be taken by the Stokes number, we begin by considering slow movement (i.e. large $\mathcal{T}$) in§ 3, and thus neglect terms of size $\mathrm{St}$. This will allow us to derive a reduced order equation (an ODE) for the deflection of the cupula that can be compared to numerical results in §4. However, in §5 we consider fast movements with finite Stokes numbers, leading to an integro-differential equation for the deflection of the cupula. Both the relative stiffness $\kappa$ and the inverse bendability $b$ are treated as independent parameters. In particular, recalling that $\kappa$ is a key parameter, capturing the relative timescales of the imposed motion and the cupular relaxation, a main objective of our analysis will be to investigate how variations in $\kappa$ give rise to different solution regimes.
@@ 3 Asymptotic solution for negligible Stokes number
$\begin{split}u(r,\theta,s,t)&=u_{0}(r,\theta,s,t)+\epsilon u_{1}(r,\theta,s,t)+\epsilon^{2}u_{2}(r,\theta,s,t)+\cdots\\ v(r,\theta,s,t)&=v_{0}(r,\theta,s,t)+\epsilon v_{1}(r,\theta,s,t)+\epsilon^{2}v_{2}(r,\theta,s,t)+\cdots\\ w(r,\theta,s,t)&=w_{0}(r,\theta,s,t)+\epsilon w_{1}(r,\theta,s,t)+\epsilon^{2}w_{2}(r,\theta,s,t)+\cdots\\ p(r,\theta,s,t)&=p_{0}(r,\theta,s,t)+\epsilon p_{1}(r,\theta,s,t)+\epsilon^{2}p_{2}(r,\theta,s,t)+\cdots\\ \eta(r,\theta,t)&=\eta_{0}(r,\theta,t)+\epsilon\eta_{1}(r,\theta,t)+\epsilon^{2}\eta_{2}(r,\theta,t)+\cdots\end{split}$ (16)
Substitution of (16) into the dimensionless problem (12) yields a system of linear equations; we shall retain terms up to and including $\mathcal{O}(\epsilon)$ since they are required to explain the symmetry breaking phenomenon that was observed numerically (see fig. 3).
@@ 3.1 Expanded solution
Substitution of the formal expansion (16) into the governing equations (12), yields the following balance after discarding terms of order $\epsilon^{2}$ and higher:
  • $\mathcal{O}(1):$
    $f(t)$ $=-\kappa\frac{\partial p_{0}}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial w_{0}}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial^{2}w_{0}}{\partial\theta^{2}},$ (17a)
@@
    $\nabla^{4}$ $\eta_{0}-b^{-1}\nabla^{2}\eta_{0}=\Delta p_{0},$ (17d)
  • $\mathcal{O}(\epsilon):$
    $f(t)\epsilon r\cos\theta$ $=-\kappa\frac{\partial p_{1}}{\partial s}+\kappa r\cos\theta\frac{\partial p_{0}}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial w_{1}}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial^{2}w_{1}}{\partial\theta^{2}}+\cos\theta\frac{\partial w_{0}}{\partial r},$ (18a)
    $\kappa\frac{\partial p_{1}}{\partial r}=\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}\cos\theta,\quad\frac{\kappa}{r}\frac{\partial p_{1}}{\partial\theta}=-\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}\sin\theta,$ (18b)
@@
    $\nabla^{4}\eta_{1}-b^{-1}\nabla^{2}\eta_{1}+\varrho r\cos\theta f(t)=\Delta p_{1}.$ (18d)
We identify the usual lubrication/boundary layer theory result that the leading order pressure is constant along a cross section (Craster & Matar, 2006). Furthermore, we note that the differential operators acting on the continuity and momentum equations are the same as in cylindrical coordinates (Batchelor, 1973).
Our starting point is the continuity equation, which can be integrated over a cross-section to deduce that the flux $Q=\int_{0}^{2\pi}\int_{0}^{a(s)}rw(r,\theta,s,t)\,\mathrm{d}r\,\mathrm{d}\theta$ is conserved in the $s$ direction, i.e. $\partial Q/\partial s=0$. This means that the flux is exclusively a function of time, a fact we will exploit to derive a reduced-order equation. Invoking symmetry we now seek an axisymmetric solution with $v_{0}=0$ and leading order terms independent of $\theta$.
Turning our attention to the $\mathcal{O}(\epsilon)$ problem, we note that $p_{1}(r,\theta,s,t)$ may be decomposed into a pressure gradient along the duct axis and an $s$-independent pressure variation due to centrifugal effects, so that
$p_{1}(r,\theta,s,t)=\bar{p}_{1}(s,t)+\frac{\dot{\Omega}_{0}\mathcal{T}^{2}F(t)^{2}}{\kappa}r\cos\theta,$ (19)
and as we only require the $s$ component of the pressure gradient in the computation of the axial velocity, we may safely ignore the $s$-independent component of $p_{1}$ and use $\bar{p}_{1}(s,t)$ in its place. Moreover, the first order pressure jump across the cupula $\Delta p_{1}=p_{1}(s=2\pi,t)-p_{1}(s=0,t)=\Delta\bar{p}_{1}$.

### v2

@@ 2.4 Scalings and non-dimensionalization
$\begin{split}&\hat{r}=ar,\quad\hat{s}=Rs,\quad\hat{z}=az,\quad\hat{t}=\mathcal{T}t,\quad\hat{w}=\frac{R\Omega_{0}a^{2}}{\mathcal{T}\nu}w,\\ &\quad\hat{p}=\frac{\mu R\mathcal{U}}{a^{2}}p,\quad\hat{\boldsymbol{u}}_{s}=\frac{a^{2}R\Omega_{0}}{\nu}\boldsymbol{u}_{s},\quad\hat{\boldsymbol{\tau}}=\frac{EaR\Omega_{0}}{\nu}\boldsymbol{\tau},\quad\hat{\Omega}=\Omega_{0}\Omega(t).\end{split}$ (8)
(‘Unhatted’ variables are therefore dimensionless counterparts of the corresponding hatted variables.) Here $\mathcal{T}$ is the timescale of variation of the forcing and $\nu=\mu/\rho$ is the kinematic viscosity of the endolymph. The velocity scale $\mathcal{U}=R\Omega_{0}a^{2}/(\mathcal{T}\nu)$ is chosen to balance the viscous forces with the Euler force. The deformation scale is chosen to balance the kinematic condition (6a). The pressure scale is chosen viscously to enforce incompressibility. Note that $\Omega(t)$ is the dimensionless angular velocity and $\dot{\Omega}(t)$ is the dimensionless angular acceleration.
@@ 2.4.1 Dimensionless equations
Following the rescaling of equations (4)–(6a), the dimensionless continuity equation (4) reads
@@
$\begin{split}\epsilon&\mathrm{St}\frac{\partial u}{\partial t}+\epsilon^{3}\mathrm{Re}\left(u\frac{\partial u}{\partial r}+\frac{v}{r}\frac{\partial u}{\partial\theta}+\frac{w}{h}\frac{\partial u}{\partial s}-\frac{v^{2}}{r}-\frac{1}{\epsilon}\frac{w^{2}}{h}\cos\theta\right)\\ &-2\Omega_{0}\mathcal{T}\mathrm{St}\Omega(t)w\cos\theta-\Omega_{0}\mathcal{T}\Omega(t)^{2}h\cos\theta\\ &=-\frac{1}{\epsilon}\frac{\partial p}{\partial r}+\frac{\epsilon}{rh}\left[\frac{r}{h}\frac{\partial}{\partial s}\left(\epsilon^{2}\frac{\partial u}{\partial s}-\frac{\partial}{\partial r}(hw)\right)-\frac{\partial}{\partial\theta}\left(\frac{h}{r}\frac{\partial}{\partial r}(rv)-\frac{h}{r}\frac{\partial u}{\partial\theta}\right)\right],\end{split}$ (10a)
$\begin{split}\epsilon&\operatorname{St}\frac{\partial v}{\partial t}+\epsilon^{2}\operatorname{Re}\left(\epsilon u\frac{\partial v}{\partial r}+\epsilon\frac{v}{r}\frac{\partial v}{\partial\theta}+\epsilon\frac{w}{h}\frac{\partial v}{\partial s}+\epsilon\frac{uv}{r}+\frac{w^{2}}{h}\sin\theta\right)\\ &+2\Omega_{0}\mathcal{T}\mathrm{St}\Omega(t)w\sin\theta+\Omega_{0}\mathcal{T}\Omega(t)^{2}h\sin\theta\\ &=-\frac{1}{\epsilon}\frac{1}{r}\frac{\partial p}{\partial\theta}+\frac{\epsilon}{h}\frac{\partial}{\partial r}\left[\frac{h}{r}\left(\frac{\partial}{\partial r}(rv)-\frac{\partial u}{\partial\theta}\right)\right]-\frac{\epsilon}{rh^{2}}\left[\frac{\partial}{\partial\theta}\left(h\frac{\partial w}{\partial s}\right)-\epsilon^{2}r\frac{\partial^{2}v}{\partial s^{2}}\right],\end{split}$ (10b)
$\begin{split}\mathrm{St}&\frac{\partial w}{\partial t}+\dot{\Omega}(t)(1+\epsilon r\cos\theta)+\epsilon 2\mathrm{St}\mathcal{T}\Omega_{0}\Omega(t)(u\cos\theta-v\sin\theta)\\ +&\epsilon^{2}\operatorname{Re}\left(u\frac{\partial w}{\partial r}+\frac{v}{r}\frac{\partial w}{\partial\theta}+\frac{w}{h}\frac{\partial w}{\partial s}+\epsilon\frac{uw}{h}\cos\theta-\epsilon\frac{vw}{h}\sin\theta\right)\\ &=-\frac{1}{h}\frac{\partial p}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left[\frac{r}{h}\left(\frac{\partial}{\partial r}(hw)-\epsilon^{2}\frac{\partial u}{\partial s}\right)\right]+\frac{1}{r^{2}}\frac{\partial}{\partial\theta}\left[\frac{1}{h}\frac{\partial}{\partial\theta}(hw)-\epsilon^{2}\frac{r}{h}\frac{\partial v}{\partial s}\right].\end{split}$ (10c)
@@
$\epsilon\frac{\rho_{s}}{\rho}\frac{1}{\kappa}\left(\mathrm{St}\frac{\partial^{2}\boldsymbol{u}_{s}}{\partial t^{2}}+\frac{\partial\mathbf{\Omega}}{\partial t}\times\mathbf{x}+2\mathrm{St}\Omega_{0}\mathcal{T}\mathbf{\Omega}\times\frac{\partial\boldsymbol{u}_{s}}{\partial t}+\Omega_{0}\mathcal{T}\mathbf{\Omega}\times(\mathbf{\Omega}\times\mathbf{x})\right)=\boldsymbol{\nabla}\cdot\boldsymbol{\tau}.$ (11)
The only inhomogeneous boundary conditions for the solid problem are
$\tau_{zz}(r,\theta,z=\pm\beta/2,t)=\pm\Delta p(r,\theta,t)/(2\kappa),$ (12)
with $\beta=t_{h}/a$ the dimensionless thickness of the cupula. The non-dimensionalization procedure introduces several dimensionless parameters. We shall see that the most important of these are the stiffness
$\kappa=\frac{E\mathcal{T}a}{R\mu},$ (13)
@@
which measures the time scale of vorticity diffusion across the channel width, $a^{2}/\nu$, to the time scale of motion, $\mathcal{T}$. (Note that this version of the Stokes number arises in Stokes’ second problem and is sometimes replaced by the Womersley number, $\mathrm{Wm}=\mathrm{St}^{1/2}$.) We also introduce the Reynolds number of the flow, $\mathrm{Re}=\rho\mathcal{U}R/\mu$, as well as the density ratio, $\rho_{s}/\rho$ and the timescale ratio $\Omega_{0}\mathcal{T}$. To understand the relative size (and hence importance) of these parameters, we next discuss characteristic parameter values and typical sizes of dimensionless parameters.
@@ 2.4.2 Parameter values
[…]
The most challenging parameter to identify is $E$. We can infer the value of the bending stiffness from the results of Selva et al. (2009), who suggest an extremely low value of the Young’s modulus, $E\sim 5$ Pa. Other authors quote simply an estimate for the bending stiffness of the cupula when modelled as a thin plate, $B=Et_{h}^{3}/(12(1-\nu_{s}^{2}))\sim 10^{-10}\mathrm{~N\,m}$ (Rabbitt & Damiano, 1992), from which a similar Pa-like Young’s modulus is inferred.
@@
In terms of the motion, usual ranges of operation for humans are $\dot{\Omega}_{0}\sim 1$ s-2 and $\mathcal{T}$ can be anywhere between 0.01 and 10 seconds. We are now in a position to make informed estimates of the sizes of the dimensionless groups appearing in (10)-(11). Both the Stokes number and the relative stiffness $\kappa$ depend on the forcing time scale $\mathcal{T}$. In particular, $\mathrm{St}\ll 1$ for $\mathcal{T}>1$ s, but the Stokes number may be large for faster movements. Similarly, $\kappa$ can acquire a large range of values. More usefully, the solid’s inertia is generally of size $\epsilon\rho_{s}\mathrm{St}/(\rho\kappa)=\rho_{s}^{2}/(E\mathcal{T})<10^{-2}$ for all physical values of $\mathcal{T}$ and hence may be safely neglected to leading order in $\epsilon$, with similar statements for all the other inertial terms in (11). We will however include this terms in our computation of the first order correction. Similarly, the reduced Reynolds number, $\epsilon^{2}\mathrm{Re}\approx 10^{-3}$ and nonlinear advection terms in (10) may be neglected. Finally, the dimensionless thickness of the cupula ranges from $\beta=0.1$ to $\beta=2$.
@@
The size of the dimensionless groups discussed above will inform our choices when neglecting terms in the governing equations (9)-(10). In particular, we will exploit the smallness of $\epsilon$ to neglect terms of order $\epsilon^{2}\mathrm{Re}$, and the smallness of $\epsilon\rho_{s}\mathrm{St}/(\rho\kappa)$ to neglect inertial terms in the Navier equation (11). The last simplification is particularly useful as it will allow us to simplify the leading order solid problem significantly to $\boldsymbol{\nabla}\cdot\boldsymbol{\tau}=\boldsymbol{0}$.
@@
We now turn to consider in more detail the behaviour of the model for the typical parameter values already discussed. Given the broad range of values that may be taken by the Stokes number, we begin by considering slow movement (i.e. large $\mathcal{T}$) in§3, and thus neglect terms of size $\mathrm{St}$. This will allow us to derive a reduced order equation (an ODE) for the deflection of the cupula that can be compared to numerical results in §4. However, in §5 we consider fast movements with finite Stokes numbers, leading to an integro-differential equation for the deflection of the cupula.
Both the relative stiffness $\kappa$ and the dimensionless cupular thickness $\beta$ are treated as independent parameters. In particular, recalling that $\kappa$ is a key parameter, capturing the relative timescales of the imposed motion and the cupular relaxation, a main objective of our analysis will be to investigate how variations in $\kappa$ give rise to different solution regimes.
@@ 3 Asymptotic solution for negligible Stokes number
$\begin{split}u(r,\theta,s,t)&=u_{0}(r,\theta,s,t)+\epsilon u_{1}(r,\theta,s,t)+\epsilon^{2}u_{2}(r,\theta,s,t)+\cdots\\ v(r,\theta,s,t)&=v_{0}(r,\theta,s,t)+\epsilon v_{1}(r,\theta,s,t)+\epsilon^{2}v_{2}(r,\theta,s,t)+\cdots\\ w(r,\theta,s,t)&=w_{0}(r,\theta,s,t)+\epsilon w_{1}(r,\theta,s,t)+\epsilon^{2}w_{2}(r,\theta,s,t)+\cdots\\ p(r,\theta,s,t)&=p_{0}(r,\theta,s,t)+\epsilon p_{1}(r,\theta,s,t)+\epsilon^{2}p_{2}(r,\theta,s,t)+\cdots\\ \boldsymbol{\tau}(r,\theta,z,t)&=\boldsymbol{\tau}_{0}(r,\theta,z,t)+\epsilon\boldsymbol{\tau}_{1}(r,\theta,z,t)+\epsilon^{2}\boldsymbol{\tau}_{2}(r,\theta,z,t)+\cdots\\ \boldsymbol{u}_{s}(r,\theta,z,t)&=\boldsymbol{u}_{s0}(r,\theta,z,t)+\epsilon\boldsymbol{u}_{s1}(r,\theta,z,t)+\epsilon^{2}\boldsymbol{u}_{s2}(r,\theta,z,t)+\cdots\end{split}$ (15)
Substitution of (15) into the dimensionless problem (10) yields a system of linear equations; we shall retain terms up to and including $\mathcal{O}(\epsilon)$ since they are required to explain the symmetry breaking phenomenon that was observed numerically (see figure 3).
@@ 3.1 Expanded solution
Substitution of the formal expansion (15) into the governing equations (10) yields the following leading order balance
$\dot{\Omega}(t)$ $=-\frac{\partial p_{0}}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial w_{0}}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial^{2}w_{0}}{\partial\theta^{2}},$ (16a)
@@
$\boldsymbol{\nabla}\cdot\boldsymbol{\tau}_{0}=\boldsymbol{0},\quad\tau_{zz0}(r,\theta,z=\pm\beta/2,t)=\pm\frac{1}{2\kappa}\Delta p_{0}(t),$ (16d)
We identify the usual lubrication/boundary layer theory result that the leading order pressure is constant along a cross section (Craster & Matar, 2006; Tavakol et al., 2017). This has the direct consequence that the pressure difference $\Delta p_{0}(t)=\int_{\epsilon\beta/2}^{2\pi-\epsilon\beta/2}\partial p_{0}/\partial s\,\mathrm{d}s$ is only a function of time. Hence, we may seek a solution to the solid problem of the form $\boldsymbol{u}_{s0}(r,\theta,z,t)=\Delta p_{0}(t)\boldsymbol{\bar{u}}_{s0}(r,z)/\kappa$, where $\boldsymbol{\bar{u}}_{s0}=(\bar{u}_{s0}(r,z),0,\bar{w}_{s0}(r,z))^{\intercal}$ is an axisymmetric solution to (11) satisfying the normalized boundary condition $\bar{\tau}_{zz0}=\pm 1/2$ at $z=\pm\beta/2$. We can thus solve the leading order solid problem independently from the fluid problem. Moreover, motivated by results from the numerical simulations indicating $w_{s0}(r,z)$ varies slowly over $z$, we introduce the depth-averaged cupular deflection $\eta(r,\theta,t)=\eta_{0}(r,t)+\epsilon\,\eta_{1}(r,\theta,t)+\cdots$ defined as
$\eta(r,\theta,t)=\frac{1}{\beta}\int_{-\beta/2}^{\beta/2}w_{s}(r,\theta,z,t)\mathrm{d}z.$ (17)
In Appendix B.2 we show that when $\mathrm{St}\ll 1$ we can decompose $\eta$ multiplicatively as
$\eta_{0}(r,\theta,t)=\frac{\Delta p_{0}(t)}{\kappa}\bar{\eta}(r,\theta),\quad\bar{\eta}(r,\theta)=\frac{1}{\beta}\int_{-\beta/2}^{\beta/2}\bar{w}_{s}(r,z,t)\mathrm{d}z,$ (18)
so that, to leading order, the cupular deflection and the pressure jump are proportional. We may obtain a polynomial approximation for $\bar{\eta}_{0}(r)$ using techniques from Barber (2010). The final solution (see Appendix B.2 for details on the calculation) is
$\bar{\eta}_{0}(r;\beta)=\frac{1}{\beta}\int_{-\beta/2}^{\beta/2}w_{s0}(r,z;\beta)\mathrm{d}z=\frac{3}{16}\frac{1-\nu_{s}^{2}}{\beta^{3}}(1-r^{2})^{2}+\frac{1}{20}\frac{(1+\nu_{s})(12-\nu_{s})}{\beta}(1-r^{2}).$ (19)
Having obtained a form of solution for the solid problem in terms of the pressure jump, we turn our attention to the fluid problem. We remark that the differential operators acting on the continuity and momentum equations are the same as in cylindrical coordinates (Batchelor, 1973), and invoking symmetry we now seek an axisymmetric solution with $v_{0}=0$ and leading order terms independent of $\theta$, with $w_{0}=w_{0}(r,s,t)$. The leading order kinematic boundary condition may be written as
$\frac{\partial\eta_{0}}{\partial t}=w_{0}(r,s=0,t)=w_{0}(r,s=2\pi,t),\quad u_{0}(r,s=0,t)=u_{0}(r,s=2\pi,t)=0.$ (20)
Our starting point is the continuity equation, which can be integrated over a cross-section to deduce that the flux $Q=\int_{0}^{2\pi}\int_{0}^{a(s)}rw(r,\theta,s,t)\,\mathrm{d}r\,\mathrm{d}\theta$ is conserved in the $s$ direction, i.e. $\partial Q/\partial s=0$. This means that the flux is exclusively a function of time, a fact we will exploit to derive a reduced-order equation. Turning our attention to the $\mathcal{O}(\epsilon)$ problem, the equations read
$\dot{\Omega}(t)r\cos\theta$ $=-\frac{\partial p_{1}}{\partial s}+r\cos\theta\frac{\partial p_{0}}{\partial s}+\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial w_{1}}{\partial r}\right)+\frac{1}{r^{2}}\frac{\partial^{2}w_{1}}{\partial\theta^{2}}+\cos\theta\frac{\partial w_{0}}{\partial r},$ (21a)
$\frac{\partial p_{1}}{\partial r}=\Omega_{0}\mathcal{T}\Omega(t)^{2}\cos\theta,\quad\frac{1}{r}\frac{\partial p_{1}}{\partial\theta}=-\Omega_{0}\mathcal{T}\Omega(t)^{2}\sin\theta,$ (21b)
@@
$\frac{\rho_{s}}{\rho}\frac{1}{\kappa}\left(\frac{\partial\boldsymbol{\Omega}}{\partial t}\times\mathbf{x}+\Omega_{0}\mathcal{T}\boldsymbol{\Omega}\times(\boldsymbol{\Omega}\times\mathbf{x})\right)=\boldsymbol{\nabla}\cdot\boldsymbol{\tau}_{1},$ (21d)
$\tau_{zz1}(r,z=\pm\beta/2,t)=\pm\Delta p_{1}(r,t)/(2\kappa),$ (21e)
$\Delta p_{1}(r,t)=\Delta p_{1}^{\text{outer}}(t)+\Delta p_{1}^{\text{BL}}(r,t).$ (21f)
Here, the pressure difference due to the outer flow is $\Delta p_{1}^{\text{outer}}(t)=\int_{\epsilon\beta/2}^{2\pi-\epsilon\beta/2}\partial p_{1}/\partial s\,\mathrm{d}s$ and $\Delta p_{1}^{\text{BL}}(r,t)$ is a contribution from the boundary layer that forms near the cupula (see Appendix F for details). We note that $p_{1}(r,\theta,s,t)$ may be decomposed into a pressure gradient along the duct axis and an $s$-independent pressure variation due to centrifugal effects, so that
$p_{1}(r,\theta,s,t)=\bar{p}_{1}(s,t)+\Omega_{0}\mathcal{T}\Omega(t)^{2}r\cos\theta,$ (22)
and as we only require the $s$ component of the pressure gradient in the computation of the axial velocity, we may safely ignore the $s$-independent component of $p_{1}$ and use $\bar{p}_{1}(s,t)$ in its place. Moreover, the first order pressure jump across the cupula is $\Delta p_{1}^{\text{outer}}=p_{1}(s=2\pi,t)-p_{1}(s=0,t)=\Delta\bar{p}_{1}^{\text{outer}}$.